OpenStudy (anonymous):

free medals even if you get the question wrong! HELP! Which is the graph of the system of equations? 2x + 3y = 12 4x + y = 8

OpenStudy (anonymous):

plz show a graph to help me understand!

OpenStudy (tmak11):

one sec

OpenStudy (anonymous):

Thank you.

OpenStudy (tmak11):

heres the graph https://www.desmos.com/calculator/ui6idbmuuv

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

one more plz Solve the following system of equations by linear combination: 3x + 2y = 7 5x – 4y = 19 A. The solution is (5, 1). B. There is no solution. C. There are an infinite number of solutions. D. The solution is (3, –1).

OpenStudy (tmak11):

5x ? 4y its all question marks

OpenStudy (anonymous):

5x-4y=19

OpenStudy (tmak11):

thanks

OpenStudy (tmak11):

give me a sec

OpenStudy (tmak11):

the lines intersect at 3,-1

OpenStudy (anonymous):

okay, so one solution.?.

OpenStudy (tmak11):

both lines are linear, meaning that there is only one possible solution

OpenStudy (anonymous):

Thank you so much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (tmak11):

now if they were greator than or less than there would be many many solutions